# Safety-guaranteed and Goal-oriented Semantic Sensing, Communication, and Control for Robotics

### 13 Mar 2026 v1

@@ IV-A Sensing
Sensing is the process of robotic systems to perceive environments, but several challenges limit the ability to ensure safety and task effectiveness. At the object level, insufficient semantic representations extraction, difficulties in point cloud segmentation, and inefficient object recognition for unknown objects and aerial views, compromise perception accuracy. At the environmental level, high computational costs in multi-sensor data processing and asynchronous multi-modal sensor data degrade situational awareness and hinder timely, safety-critical decisions. The following subsections detail these challenges and potential GSC research directions.
C1: Insufficient Semantic Representations Extraction. Robots need to extract information from sensor data to recognize their surrounding objects. However, current models only extract limited information (e.g., object type), resulting in inefficient object recognition, compromising both safety and effectiveness. For instance, in real-time path planning, ignoring critical information like relative distance between robots and obstacles can lead to unsafe control decisions, violating safety distance, thereby increasing collision risk and mapping time.
S1: Semantic Representations Definition. To address this challenge, we define semantic representations that quantify the importance for task safety and effectiveness. As safety requirements and effectiveness metrics vary by tasks, these representations are tailored to each context. Table III summarizes them across three given use cases, highlighting how they differ based on task-specific requirements.
@@
S2: Semantic-based Segmentation. To solve this, the state-of-the-art approach uses contrastive learning to sub-sample and group same-class points near each boundary, followed by average pooling and class distribution analysis for multi-scale refinement. However, this is computationally expensive, as all objects are treated equally regardless of their importance to the task. A more efficient semantic-based approach could be proposed, which detects important objects and prioritizes boundary segmentation for them, reducing computation while preserving safety and effectiveness.
@@
C3: Inefficient Object Recognition. Current object recognition models are inefficient, particularly for unknown objects or those imaged from varying altitudes and angles (e.g., top-down), which may cause distortion and misclassifications. Such errors can violate safety requirements on grasping force, leading to slippage and increased execution time. To address these, we propose the following solutions.
@@
S3.1: Semantic-based Unknown Object Recognition. Existing methods for unknown object recognition largely adopt open-set detection, using uncertainty estimation or knowledge transfer to distinguish known from unknown classes. While effective for classification, it cannot ensure safe, reliable task execution. Thus, a semantic-based approach could be proposed, emphasizing semantic representations (e.g., texture, material). This can also be extended to unknown-object grasping, where current methods struggle with generalization and sim-to-real transfer. Semantic data augmentation can enhance policy robustness, while domain adaptation and feature fusion can improve alignment between simulated and real-world distributions.
@@
S3.2: Multi-view Image Fusion. UAV aerial images are typically captured from bird’s-eye views, rendering side-view–oriented 3D object detection models ineffective. A common solution is monocular 3D detection with height offsets, but it is highly sensitive to height errors and viewpoint distortions. Multi-view image fusion could offer a robust solution by capturing objects from multiple angles and enforcing: (1) projection consistency, aligning 2D projections of predicted 3D boxes with ground truth; and (2) multi-view consistency, ensuring stable 3D boxes across views. This reduces reliance on height estimates and mitigates viewpoint distortion.
@@
C4: Multiple Sensor Data Processing. Robots transmit large volumes of sensor data for edge processing, but limited computational resources can cause delays that compromise safety and effectiveness. In multi-robot exploration, local images are queued for processing. Delays in critical images’ (e.g., robots nearing obstacles) processing may violate safety distance requirements, leading to collisions and longer mapping times. To address this, we propose several potential solutions.
@@
S4.1: Semantic-based Queue Priority Design. Traditional queuing methods, such as first-in–first-out (FIFO) and first-in–last-out, prioritize data solely by arrival time, which can delay processing of critical sensor data. For example, in FIFO, late-arriving but safety-critical data may be processed too late. Thus, a semantic-based queue priority scheme could be proposed to rank data by its importance. Specifically, more important sensor data (e.g., images with numerous obstacles) is prioritized for earlier processing.
@@
S4.2: Semantic-based Task Offloading. Offloading sensor data processing from edge to cloud can reduce computational delays, but existing methods rely solely on bit-level metrics, ignoring semantic representations. This may lead to inefficient offloading, such as processing non-critical images (e.g., static robots) while delaying critical ones (e.g., robots near obstacles), increasing collision risk. Thus, a semantic-aware offloading approach could be proposed to offload sensor data processing based on its importance to safety and effectiveness.
@@
S5: Multi-sensor Data Synchronization. Existing synchronization methods include software-based schemes, such as network time protocol (NTP), and hardware-based schemes, such as IEEE 1588-2008 precision time protocol (PTP). While both can achieve sub-microsecond precision, NTP suffers from limited accuracy in dynamic networks, and PTP requires specialized hardware, limiting deployment. As not all data demands equal precision, a hierarchical framework could be proposed to allocate higher precision to safety/effectiveness-critical data (e.g., tactile signals) and lower precision to others.
@@ IV-B Communication
Communication refers to the process of transmitting both sensor and C&C data, with the following challenges and potential research directions.
@@
C6: Limited Communication Resources. Limited communication resources for uplink sensor and downlink C&C data transmission could cause delays/distortions that compromise safety and efficiency, especially for dense point cloud data (e.g., Ford 01 dataset: 120 million points over 1500 frames at 54 bits/point). This may prevent timely obstacle recognition, increasing collision risks and mapping time. To address this, we propose the following potential research directions.
@@
S6.1: Semantic-based Sensor Data Compression. One approach is to reduce sensor data size via compression. However, existing methods focus on full data reconstruction rather than selectively preserving semantic representations, leading to inefficiency. Thus, we proposed the following potential solutions:
@@
  • Semantic-based Traditional Compression Redesign: Traditional compression applies uniform strategies across data, yielding consistent ratios regardless of each part’s importance to safety and effectiveness. Thus, a semantic-based redesign could be proposed to guide compression by importance. Specifically, in transform-based methods (e.g., JPEG), apply finer quantization to critical blocks. In octree-based point-cloud compression, use deeper subdivisions for key sub-cubes. In graph-based compression, retain high semantic ranked nodes while discarding less significant ones.
@@
  • Semantic-based Rate-distortion Optimization: Learned compression methods typically optimize bit-level rate–distortion functions, which may be too stringent for safe/effective task execution and lead to suboptimal compression. For example, a method that retains only target objects while discarding the background may enable effective grasping but is penalized under conventional rate–distortion metrics due to excessive image distortion. To address this, semantic-based rate–distortion functions can replace bit rate with semantic representation rate (semantic content transmitted per time slot) and distortion with semantic representation distortion (deviation in reconstructed semantics).
@@
S6.2: Semantic-aware Resource Allocation and Scheduling. Another research direction is to jointly optimize resource allocation and scheduling for sensor or C&C data transmission. Existing schemes, such as proportional fair scheduling and frequency-domain packet scheduling (FDPS), focus solely on bit-level uplink performance, resulting in inefficient resource utilization. For example, FDPS may allocate minimal resources to critical data (e.g., images containing target objects) under poor channel conditions, delaying transmission.This highlights the need for semantic-aware resource allocation and scheduling, where more communication resources are allocated to more important sensor or C&C data to ensure timely and reliable transmission.
@@
S6.3: Semantic-aware Dynamic Transmission. Existing fixed-period transmission of sensor or C&C data wastes resources on low-importance data. This brings the need to design a semantic-aware dynamic transmission approach, which adapts transmission rate based on data importance. Specifically, only essential data is transmitted, with rates increased in high-risk scenarios (e.g., dense obstacles) or under more stringent safety requirements, enabling more frequent delivery of critical data.
@@ IV-C Control
Control involves generating high-level decisions and executing low-level commands, but several challenges hinder its ability to guarantee safety and task effectiveness. At the higher level, complex coordinated decision-making and unstable real-time path replanning can result in unsafe or ineffective actions. At the lower level, incorrect C&C data may control the robot into unsafe states. The following subsections outline these challenges and potential research directions.
C7: Coordinated Decision-making. Effective multi-robot collaboration requires real-time coordinated decision-making, which becomes more challenging as robot numbers grows. This highlights the need for a coordinated decision-making strategy that dynamically adapts to task progress and inter-robot dependencies, ensuring both safety and effectiveness.
S7: Multi-robot Collaboration. Two common coordinated decision-making strategies are game theory and cooperative MADRL. Game theory yields stable but suboptimal strategies by balancing each robot’s benefits (e.g., area explored), costs (e.g., travel distance), and safety (e.g., collision risk). Cooperative MADRL maximizes shared global performance but potentially inducing unsafe individual actions. A promising approach is to combine decentralized large language models (LLMs) with MPC. Each robot uses an LLM to exchange capabilities and intentions, collaboratively generating a high-level cooperation plan. MPC then translates this plan into low-level motion commands and validates safety via future-state simulations before execution.
@@
S8: Fast and Adaptive Replanning. UAVs and mobile robots often employ DRL-based replanners to decide when to replan, but lack dynamic control over replanning frequency. Vision-based methods with vision–language models are commonly used in robot arms to infer spatial relationships (e.g., obstacle proximity), but slow processing can yield outdated decisions. Thus, a promising unified solution is to design a hierarchical replanning framework: 1) a fast, low-level replanner ensures real-time safety via short-horizon control, triggered when safety is at risk; and 2) a slower, high-level replanner refines the global trajectory to optimize long-term effectiveness.
C9: C&C Data Execution. Safe and effective task execution requires robots to follow optimal C&C data, as wrong C&C data can hinder the robot’s ability to maintain safety distance requirement, risking damage to the robot. This highlights the need to ensure optimal C&C execution.
S9: Semantic-based C&C Data Execution. Similar to sensor data, incoming C&C data is queued for execution in robots. Existing methods execute only the latest data, which can be unsafe if it is delayed or outdated. A semantic-based execution strategy could be potentially proposed to prioritize C&C data by its importance to task safety and effectiveness. The importance could be ranked by freshness, quantified via age of information, and data value, measured by value of information.
@@ V Case Study
To validate the effectiveness of our proposed safety-guaranteed and goal-oriented semantic communication solution, we implement the semantic-based C&C data execution solution, namely SemCE, in a UAV target tracking task. In this setup, a UAV is controlled to track a moving target over a total of $T$ time slots. At time slot $t\in\{1,2,\ldots,T\}$, let $d_{t}$ denotes the Euclidean distance between the UAV and the target. Let $d$, $d_{\mathrm{s}}$ and $d_{\mathrm{e}}$ denotes the safety distance requirement and tracking threshold, with $d_{\mathrm{e}}>d_{\mathrm{s}}>0$. The tracking status at time slot $t$ is classified as: 1) unsafe if $d_{t}\leq d_{\mathrm{s}}$; 2) successful if $d_{\mathrm{s}}<d_{t}\leq d_{\mathrm{e}}$; 3) unsuccessful if $d_{t}>d_{\mathrm{e}}$. Effectiveness and safety performance are measured by the tracking success and safety rate. The tracking success rate refers to the proportion of successful time slots, while the safety rate is defined as $\frac{T_{\mathrm{s}}}{T}$, where $T_{\mathrm{s}}$ is the number of time slots in which tracking is safe. C&C data is generated using a PID controller (proportional gain = 0.5).
@@
Fig. 2 compares the safety rate (blue bars) and tracking success rate (orange bars) of the baseline and SemCE under varying tracking thresholds $d_{\mathrm{e}}$ and safety distance requirements $d_{\mathrm{s}}$. It is noticed that SemCE consistently outperforms the baseline in both metrics. In Fig. 2 (a), with $d_{\mathrm{s}}$ fixed at 1 m and varying $d_{\mathrm{e}}$, both schemes maintain high safety rates as the safety requirement is not stringent. As $d_{\mathrm{e}}$ decreases and approaches $d_{\mathrm{s}}$, the success rate of both schemes declines due to a narrower tracking range. In Fig. 2 (b), with $d_{\mathrm{e}}$ fixed at 5 m and increasing $d_{\mathrm{s}}$, both safety and success rates decline due to more stringent safety requirements. Notably, SemCE exhibits a much smaller performance drop. These results validate that our proposed safety-guaranteed goal-oriented semantic communication solution significantly enhances both safety and effectiveness.

### 26 Apr 2026 v2

@@ IV-A Sensing
Sensing is the process of robotic systems to perceive their environment, but several challenges limit the ability to ensure safety and task effectiveness. At the object level, insufficient semantic representation extraction, difficulties in point cloud segmentation, and inefficient object recognition for unknown objects and aerial views, compromise perception accuracy. At the environmental level, high computational costs in multi-sensor data processing and asynchronous multi-modal sensor data degrade situational awareness and hinder timely, safety-critical decisions. The following subsections detail these challenges and potential research directions.
C1: Insufficient Semantic Representations Extraction. Robots need to extract information from sensor data to recognize their surrounding objects. However, current models only extract limited information (e.g., object type), resulting in inefficient object recognition, compromising both safety and effectiveness. For instance, in real-time path planning, ignoring critical information like relative distance between robots and obstacles can lead to unsafe control decisions, violating the safety distance, thereby increasing collision risk and mapping time.
D1: Semantic Representations Definition. To address this challenge, we define semantic representations that quantify the importance for task safety and effectiveness. As safety requirements and effectiveness metrics vary by tasks, these representations are tailored to each context. Table III summarizes them across the three given use cases, highlighting how they differ based on task-specific requirements.
@@
D2: Semantic-based Segmentation. To solve this, the state-of-the-art approach uses contrastive learning to sub-sample and group same-class points near each boundary, followed by average pooling and class distribution analysis for multi-scale refinement. However, this is computationally expensive, as all objects are treated equally regardless of their importance to the task. A promising future direction is to propose a more efficient semantic-based approach, which detects important objects and prioritizes boundary segmentation for them, reducing computation while preserving safety and effectiveness.
@@
C3: Inefficient Object Recognition. Current object recognition models are inefficient, particularly for unknown objects or those imaged from varying altitudes and angles (e.g., top-down), which may cause distortion and misclassifications. Such errors can violate safety requirements on grasping force, leading to slippage and increased execution time. To address these, we outline the following future research directions.
@@
D3.1: Semantic-based Unknown Object Recognition. Existing methods for unknown object recognition largely adopt open-set detection, using uncertainty estimation or knowledge transfer to distinguish known from unknown classes. While effective for classification, they cannot ensure safe, reliable task execution. Thus, One promising direction is to emphasize semantic representations (e.g., texture, material). This can also be extended to unknown-object grasping, where current methods struggle with generalization and sim-to-real transfer. Semantic data augmentation can enhance policy robustness, while domain adaptation and feature fusion can improve alignment between simulated and real-world distributions.
@@
D3.2: Multi-view Image Fusion. UAV aerial images are typically captured from bird’s-eye views, rendering side-view–oriented 3D object detection models ineffective. A common solution is monocular 3D detection with height offsets, but it is highly sensitive to height errors and viewpoint distortions. Multi-view image fusion could serve as a potential direction by capturing objects from multiple angles and enforcing: (1) projection consistency, aligning 2D projections of predicted 3D boxes with ground truth; and (2) multi-view consistency, ensuring stable 3D boxes across views. This could reduce reliance on height estimates and mitigates viewpoint distortion.
@@
C4: Multiple Sensor Data Processing. Robots transmit large volumes of sensor data for edge processing, but limited computational resources can cause delays that compromise safety and effectiveness. In multi-robot exploration, local images are queued for processing. Delays in critical images’ (e.g., robots nearing obstacles) processing may violate safety distance requirements, leading to collisions and longer mapping times. To address this, we propose several research directions.
@@
D4.1: Semantic-based Queue Priority Design. Traditional queuing methods, such as first-in–first-out (FIFO) and first-in–last-out, prioritize data solely by arrival time, which can delay processing of critical sensor data. For example, in FIFO, late-arriving but safety-critical data may be processed too late. Thus, a promising solution is to design a semantic-based queue priority scheme, which ranks data by its importance. Specifically, more important sensor data (e.g., images with numerous obstacles) is prioritized for earlier processing.
@@
D4.2: Semantic-based Task Offloading. Offloading sensor data processing from edge to cloud can reduce computational delays, but existing methods rely solely on bit-level metrics, ignoring semantic representations. This may lead to inefficient offloading, such as processing non-critical images (e.g., static robots) while delaying critical ones (e.g., robots near obstacles), increasing collision risk. Thus, one direction is to design a semantic-aware offloading approach that offloads sensor data processing based on its importance to safety and effectiveness.
@@
D5: Multi-sensor Data Synchronization. Existing synchronization methods include software-based schemes, such as network time protocol (NTP), and hardware-based schemes, such as IEEE 1588-2008 precision time protocol (PTP). While both can achieve sub-microsecond precision, NTP suffers from limited accuracy in dynamic networks, and PTP requires specialized hardware, limiting deployment. As not all data demands equal precision, one direction is to propose a hierarchical framework to allocate higher precision to safety/effectiveness-critical data (e.g., tactile signals) and lower precision to others.
@@ IV-B Communication
Communication refers to the process of transmitting both sensor and C&C packets, with the following challenges and potential research directions.
@@
C6: Limited Communication Resources. Limited communication resources for uplink sensor and downlink C&C packets transmission could cause delays/distortions that compromise safety and efficiency, especially for dense point cloud data (e.g., Ford 01 dataset: 120 million points over 1500 frames at 54 bits/point). This may prevent timely obstacle recognition, increasing collision risks and mapping time. To address this, we outline the following potential research directions.
@@
D6.1: Semantic-based Sensor Data Compression. One approach is to reduce sensor data size via compression. However, existing methods focus on full data reconstruction rather than selectively preserving semantic representations, leading to inefficiency. Thus, we outline the following potential research directions:
@@
  • Semantic-based Traditional Compression Redesign: Traditional compression applies uniform strategies across data, yielding consistent ratios regardless of each part’s importance to safety and effectiveness. Thus, a semantic-based redesign could be outlined as a direction, which guides compression by importance. Specifically, in transform-based compression (e.g., JPEG), semantic importance can guide distortion weighting and bit allocation across blocks or coefficients. In octree-based point-cloud compression, partitioning depth and point retention can be adaptively adjusted according to the semantic importance of local regions. In graph-based compression, node or edge preservation can be prioritized based on their contribution to downstream tasks.
@@
  • Semantic-based Rate-distortion Optimization: Learned compression methods typically optimize bit-level rate–distortion functions, which may be too stringent for safe/effective task execution and lead to suboptimal compression. For example, a method that retains only target objects while discarding the background may enable effective grasping but is penalized under conventional rate–distortion metrics due to excessive image distortion. To address this, one direction is to design semantic-based rate–distortion functions, which can replace bit rate with semantic representation rate (semantic content transmitted per time slot) and distortion with semantic representation distortion (deviation in reconstructed semantics).
@@
D6.2: Semantic-aware Resource Allocation and Scheduling. Another research direction is to jointly optimize resource allocation and scheduling for sensor or C&C packets transmission. Existing schemes, such as proportional fair scheduling and frequency-domain packet scheduling (FDPS), focus solely on bit-level uplink performance, resulting in inefficient resource utilization. For example, FDPS may allocate minimal resources to critical data (e.g., images containing target objects) under poor channel conditions, delaying transmission. This highlights the need for semantic-aware resource allocation and scheduling, where more communication resources are allocated to more important sensor or C&C packets to ensure timely and reliable transmission.
@@
D6.3: Semantic-aware Dynamic Transmission. Existing fixed-period transmission of sensor or C&C packets wastes resources on low-importance data. This introduces another direction to design a semantic-aware dynamic transmission approach, which adapts transmission rate based on data importance. Specifically, only essential data is transmitted, with rates increased in high-risk scenarios (e.g., dense obstacles) or under more stringent safety requirements, enabling more frequent delivery of critical data.
@@ IV-C Control
Control involves generating high-level decisions and executing low-level commands, but several challenges hinder its ability to guarantee safety and task effectiveness. At the higher level, complex coordinated decision-making and unstable real-time path replanning can result in unsafe or ineffective actions. At the lower level, incorrect C&C packets may control the robot into unsafe states. The following subsections outline these challenges and potential research directions.
C7: Coordinated Decision-making. Effective multi-robot collaboration requires real-time coordinated decision-making, which becomes more challenging as robot number grows. This highlights the need for a coordinated decision-making strategy that dynamically adapts to task progress and inter-robot dependencies, ensuring both safety and effectiveness.
D7: Multi-robot Collaboration. Two common coordinated decision-making strategies are game theory and cooperative MADRL. Game theory yields stable but suboptimal strategies by balancing each robot’s benefits (e.g., area explored), costs (e.g., travel distance), and safety (e.g., collision risk). Cooperative MADRL maximizes shared global performance but potentially inducing unsafe individual actions. A promising direction is to combine decentralized large language models (LLMs) with MPC. Each robot uses an LLM to exchange capabilities and intentions, collaboratively generating a high-level cooperation plan. MPC then translates this plan into low-level motion commands and validates safety via future-state simulations before execution.
@@
D8: Fast and Adaptive Replanning. UAVs and mobile robots often employ DRL-based replanners to decide when to replan, but lack dynamic control over replanning frequency. Vision-based methods with vision–language models are commonly used in robot arms to infer spatial relationships (e.g., obstacle proximity), but slow processing can yield outdated decisions. Thus, a promising direction is to design a hierarchical replanning framework: 1) a fast, low-level replanner ensures real-time safety via short-horizon control, triggered when safety is at risk; and 2) a slower, high-level replanner refines the global trajectory to optimize long-term effectiveness.
C9: C&C Packet Execution. Safe and effective task execution requires robots to receive correct C&C packets, as erroneous C&C packets can hinder the robot’s ability to maintain the safety distance requirement, risking damage to the robot. This highlights the need to ensure optimal C&C execution.
D9: Semantic-based C&C Packet Execution. Similar to sensor data, incoming C&C packets are queued for execution in robots. Existing methods execute only the latest data, which can be unsafe if it is delayed or outdated. Thus, one direction is to design a semantic-based execution strategy that could prioritize C&C packets by their importance to task safety and effectiveness. The importance could be ranked by freshness, quantified via age of information, and data value, measured by value of information.
@@ V Case Study
To validate the effectiveness of the proposed SA-GS directions, we implement one representative direction, namely semantic-based C&C packet execution (D9), referred to as SemCE, in a single-UAV target-tracking task. Specifically, an edge server periodically transmits C&C data to control a UAV to track a mobile target (eg., vehicle) moving over $T$ time slots. The mobile target moves along a randomly generated trajectory, with a constant speed within each time slot that may vary across time slots. We assume the mobile target can send its real-time positions to the edge server correctly all the time without error. Meanwhile, the UAV’s onboard sensor captures its real-time positions, also forwarding them to the edge server. At each time slot $t\in\{1,2,\ldots,T\}$, the edge server generates a C&C packet according to the current tracking state using a proportional integral derivative (PID) controller with proportional gain 0.5. The C&C packet is transmitted through a Rayleigh-fading channel. As a baseline, the UAV executes the most recently received C&C packet, referred to as the latest-arrived execution policy. By contrast, SemCE prioritizes received packets based on their freshness and value of information. Let $d_{t}$ denote the Euclidean distance between the UAV and the target. Let $d_{\mathrm{s}}$ and $d_{\mathrm{e}}$ denote the safety distance and tracking threshold, respectively, where $d_{\mathrm{e}}>d_{\mathrm{s}}>0$. Then, the tracking status at time slot $t$ is classified as unsafe if $d_{t}\leq d_{\mathrm{s}}$, successful if $d_{\mathrm{s}}<d_{t}\leq d_{\mathrm{e}}$, and unsuccessful if $d_{t}>d_{\mathrm{e}}$. Accordingly, safety and effectiveness are evaluated by the safety rate and tracking success rate, respectively. The safety rate is defined as $T_{\mathrm{s}}/T$, where $T_{\mathrm{s}}$ is the number of safe time slots, while the tracking success rate is the proportion of time slots satisfying $d_{\mathrm{s}}<d_{t}\leq d_{\mathrm{e}}$.
@@
Fig. 2 compares the safety rate (blue bars) and tracking success rate (orange bars) of the baseline and SemCE under two threshold settings. In Fig. 2 (a), the safety distance is fixed at $d_{\mathrm{s}}=1$ m, while the tracking threshold varies as $d_{\mathrm{e}}\in\{1.5,2.5,3.5,4.5\}$ m. In Fig. 2 (b), the tracking threshold is fixed at $d_{\mathrm{e}}=5$ m, while the safety distance varies as $d_{\mathrm{s}}\in\{1,2,3,4\}$ m. It is observed that SemCE consistently outperforms the latest-arrived baseline in both safety rate and tracking success rate across all evaluated settings. In Fig. 2 (a), with $d_{\mathrm{s}}$ fixed at 1 m and varying $d_{\mathrm{e}}$, both schemes maintain high safety rates as the safety requirement is not stringent. As $d_{\mathrm{e}}$ decreases and approaches $d_{\mathrm{s}}$, the success rate of both schemes declines due to a narrower tracking range. In Fig. 2 (b), as $d_{\mathrm{s}}$ increases with fixed $d_{\mathrm{e}}=5$ m, both safety and tracking become more challenging, leading to performance degradation for both schemes. Nevertheless, SemCE exhibits a markedly smaller drop, indicating that prioritizing fresher and more important C&C packets improves both safe operation and tracking effectiveness.